Calculate the pH of each of the following solution: 0.45MNaOCl
Answer #1

NaOCl is a strong salt : NaOCl >> Na+ + OCl-

OCl- + H2O <-----> HClO + OH-

Given that 0.45M NaOCl

K b = Kw/Ka = 3.3 x 10^-7

K b= [HClO ][ OH-] /[ OCl- ]

K b = x^2 / 0.45-x

x = [OH-]= 0.00039 M

pOH = - log oh-

pOH = 3.4

pH = 14 - 3.4 = 10.6